OpenStudy (anonymous):

Find the angle between v =3i-2j and w=-i+4j. Express your answer in degrees and round to the nearest tenth.

OpenStudy (jdoe0001):

\(\bf v =3i-2j\implies v=<3,-2> \qquad w=-i+4j\implies w=<-1,4>\\ \quad \\ \textit{angle between them will be at }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

what answer did u get. I wanna make sure i did it right